# Disentangled Source-Free Personalization for Facial Expression Recognition with Neutral Target Data

### v1

@@ Abstract
Facial Expression Recognition (FER) from videos is a crucial task in various application areas, such as human-computer interaction and health monitoring (e.g., pain, depression, fatigue, and stress). Beyond the challenges of recognizing subtle emotional or health states, the effectiveness of deep FER models is often hindered by the considerable variability of expressions among subjects. Source-free domain adaptation (SFDA) methods are employed to adapt a pre-trained source model using only unlabeled target domain data, thereby avoiding data privacy and storage issues. Typically, SFDA methods adapt to a target domain dataset corresponding to an entire population and assume it includes data from all recognition classes. However, collecting such comprehensive target data can be difficult or even impossible for FER in healthcare applications. In many real-world scenarios, it may be feasible to collect a short neutral control video (displaying only neutral expressions) for target subjects before deployment. These videos can be used to adapt a model to better handle the variability of expressions among subjects. This paper introduces the Disentangled Source-Free Domain Adaptation (DSFDA) method to address the SFDA challenge posed by missing target expression data. DSFDA leverages data from a neutral target control video for end-to-end generation and adaptation of target data with missing non-neutral data. Our method learns to disentangle features related to expressions and identity while generating the missing non-neutral target data, thereby enhancing model accuracy. Additionally, our self-supervision strategy improves model adaptation by reconstructing target images that maintain the same identity and source expression. Experimental results [^1] Our code is provided in suppl. materials and will be made public. on the challenging BioVid and UNBC-McMaster pain datasets indicate that our DSFDA approach can outperform state-of-the-art adaptation methods. Our code is available at: github.com/MasoumehSharafi/DSFDA-for-Pain-Assessment.
Keywords: Disentangled Learning, Domain Adaptation, Pain Estimation, Emotion Recognition
Figure 1: Illustration of SFDA settings using pain estimation as an example. (a) A standard subject-based SFDA setting where the source pre-trained model $F_{S}$ is adapted through a pseudo-labeling approach using unlabeled target video data from all classes (neutral and pain levels). (b) Our proposed subject-based disentangled SFDA (DSFDA) setting with neutral control target data. The generator $G$ is employed to synthesize the missing classes in the target domain based on the neutral control video. (c) Comparison of accuracy on BioVid data [46] of a source-only pre-trained Inception-V3 CNN ($F_{s}$ before adaptation), after adaptation with only neutral control data, and with our DSFDA method on neutral and generated pain data. (Locks denote that labeled source data is not available during adaptation.)
1 Introduction
@@
Facial expression recognition (FER) has attracted much attention in computer vision due to the breadth of its applications, such as pain assessment, stress estimation, and driver fatigue monitoring. Although state-of-the-art (SOTA) deep FER models [33, 40, 41, 59] have made significant progress, developing applications and models capable of characterizing expression remains a challenge. Deep learning models [31], [32], [50] have shown promising performance on various FER tasks. However, their performance can degrade significantly when there is a shift in data distributions between the source (lab) and target (operational) domains. Various unsupervised adaptation (UDA) methods [25, 29, 56, 61] have been proposed to address the distribution shifts between the source and the target domain. For instance, the work by Zeeshan et al. [56] introduces a subject-based multi-source domain adaptation (MSDA) method specifically tailored for FER tasks, where each subject is considered as an individual domain for fine-grained adaptation to the variability of expressions among subjects.
@@
State-of-the-art UDA methods [5, 19, 25, 26, 61] rely on the availability of the labeled source data during adaptation, which is impractical in real-world scenarios due to concerns for data privacy, data storage, transmission cost, and computation burden. Therefore, to reduce the reliance on source data during the adaptation process, several source-free (unsupervised) domain adaptation (SFDA) methods have been proposed [11, 24, 28, 54, 57].
@@
SFDA methods typically focus on the scenario where unlabeled target data from all classes is available for the adaptation process (see Fig. 1(a)). However, in real-world video-based health applications, person-specific data representing non-neutral expressions is typically costly or unavailable. A short neutral control video may, however, be collected for target individuals and used to personalize a model to the variability of an individual’s diverse expressions. In practice, collecting neutral target data for adaptation is generally easier than gathering non-neutral emotional data. Annotating neutral expressions is more straightforward and less subjective than interpreting emotional expressions, which can vary considerably among expert annotators. Creating scenarios to elicit specific emotions is complex, time-consuming, and raises ethical and privacy concerns. In contrast, capturing neutral expressions is less intrusive and may encourage more willing participation, as it avoids the need to act out emotional states.
@@
In this paper, a new personalized SFDA method is proposed for FER with our new setting (see Fig. 1(b)). The Disentangled SFDA (DSFDA) method is proposed to address the challenges of personalized SFDA using neutral control target data by generating the missing non-neutral expressive image data for target subjects. Using a dual-generator architecture pre-trained on source subjects, a neutral image is transformed into one that exhibits the missing non-neutral expression. By combining features from two input images, the method synthesizes expressive images that belong to the missing non-neural target expressions, enabling accurate FER predictions for target subjects even without explicit data for those expressions. Our approach is further enhanced using self-supervised learning, which ensures the reconstruction of input target images with the same identity and source expression. This process improves the model’s adaptability and prediction accuracy in the absence of source data. As shown in Fig. 1(c), DSFDA overcomes the challenges of domain shifts and incomplete target data, ensuring a high level of performance across diverse target domains.
Our contributions can be summarized as follows. (1) A novel SFDA setting is introduced, where all classes in the source data are utilized to pre-train the source model. For target adaptation, however, only neutral control target data is available to adapt the model for a target subject. (2) We propose the DSFDA method, a one-stage approach tailored to our novel setting, that combines data generation and target domain adaptation to simultaneously generate non-neutral target image data and adapt a FER model. (3) An extensive set of experimental results indicating that the FER model adapted using our DSFDA method outperforms state-of-the-art SFDA methods on the BioVid and UNBC-McMaster datasets. The method scales well to the number of subjects and is robust across various subjects, target domains, and datasets.
2 Related Work
2.1 Source-Free Domain Adaptation
@@
Model-based methods: Conti et al. [8] introduced a cluster purity score to obtain reliable pseudo-labels using $k$-means clustering. The target feature extractor is then trained using two augmentations of input images instead of raw input data. Then, the reduced target dataset is trained using the obtained reliable pseudo-labels with the self-supervised feature extractor. Guo et al. [13] incorporate Tsallis entropy into the information maximization constraint to enhance the confidence of prediction labels while reducing confidence bias. Moreover, the method utilizes virtual adversarial training by employing perturbed data alongside original target domain data to improve the model’s domain generalization ability.
@@
Unlike the promising results of SFDA methods in benchmark classification tasks, current SFDA approaches are typically designed under the assumption of an equal number of classes in both the source and target domains. However, in FER tasks, non-neutral expression data (missing classes) may be unavailable due to the high costs of data collection and annotation, with only a short neutral video available for adaptation. This leads to a label shift between the target adaptation and testing phases, in addition to the domain shift between the source and target domains, ultimately degrading classification performance. To mitigate the shift between the source and target domains, generative models are utilized to synthesize the missing classes within the target domain, thereby enriching the diversity of the target data. These models aim to produce high-quality images or videos that accurately represent the emotional expressions absent in the original target dataset. In the literature, a variety of techniques have been explored for this purpose, ranging from traditional generative adversarial networks (GANs) and variational autoencoders (VAEs) to more advanced frameworks such as conditional GANs, diffusion models, and disentanglement methods.
2.2 Facial Expression Generation and Disentanglement
@@
Generative modeling, a branch of machine learning (ML), concentrates on understanding the underlying distribution of training data to enable the generation of new samples that mimic the statistical features of that data. Over time, significant progress has been made in this field. Most common techniques like GANs [2, 3, 4, 7, 58], VAEs [22, 37, 44], and diffusion models [9, 21, 34, 38] have significantly improved the quality and variety of the samples generated. For facial expression generation, [42] developed FExGAN-Meta for meta-humans, while ImaGINator [48] employed CGANs for creating facial expression video sequences. Niinuma et al. [35] used a GAN-based approach to synthesize facial expressions for AU intensity estimation, improving performance on imbalanced and domain-shifted datasets through personalized networks. Recent works have emphasized domain generalization and adaptation for cross-corpus facial expression recognition. Yin et al. [55] introduced the FATE model, which uses contrastive learning and facial animation warping for superior cross-domain emotion recognition. Furthermore, DeepFN [15] proposed deep face normalization for robust action unit recognition across individuals. Kim et al. [53] leveraged single-image temporal cues for action unit detection, enhancing the understanding of subtle expressions.
Disentanglement in ML refers to the process of learning representations where different factors of variation in the data are isolated into distinct components [16]. It is helpful for tasks like data augmentation, image synthesis, and improving ML model performance. In the literature, numerous methods of disentanglement are used for different tasks. In their study, [47] suggest classifying these techniques according to several criteria. Disentanglement holds significant relevance for facial expression recognition, where expression features often exhibit a high correlation with the identities of subjects. Popularized by [6], GANs are frequently employed in FER disentanglement studies to effectively separate expression features from identity features, thereby improving the performance and interpretability of models [1, 50]. Additionally, disentangling other factors is possible as well. The IPD-FER method utilizes three encoders to disentangle identity, pose, and expression factors [20]. Furthermore, Halawa et al. [14] propose a GAN-based disentanglement approach that achieves the separation of expression features without relying on identity labels.
@@
Beyond GAN, other techniques are also explored for disentanglement, including the VAEs methods, which is another popular method for disentanglement [60]. Moreover, Li et al. [27] propose to use the Hilbert-Schmidt independence criterion to disentangle expression from identity features, demonstrating the application of mathematical criteria in disentanglement tasks. In the same way, mutual information minimization is a strategy for achieving this objective [39]. These studies highlight diverse methodologies aimed at achieving robust and interpretable disentangled representations in facial expression recognition.
Given only neutral control videos in the target domain, we generate video frames that represent non-neutral expressions. This two-stage approach converts the SFDA problem with only neutral target videos into a standard SFDA scenario where both domains share the same classes. First, a generator (like VAEs, GANs, or DMs) is trained on the source domain to create images with non-neutral expressions, while a source model is pre-trained on the source data. Then, during inference, the generator produces images or videos for the target data, which initially lacks these expressions. These synthetic images are combined with the target’s neutral images to adapt the source model to the target domain, improving expression recognition. However, current generative methods often struggle to capture subtle pain intensity levels due to the complexity of nuanced expressions.
SOTA methods for SFDA assume the same classes for target adaptation and testing across all settings, such as closed-set, partial-set, and open-set. However, in FER tasks, non-neutral expression data may not be available due to the challenges and costs of data collection and annotation, leaving only neutral control target video for adaptation. This results in a label shift between the target adaptation and testing phases, in addition to the domain shift between the source and target domains, ultimately degrading classification performance. While generative models are used to produce non-neutral expressions for target data, they often struggle with subtle expressions, such as varying levels of pain intensity. This limitation arises because these deep FER models find it challenging to handle the nuances required for accurate expression generation. As described in the next section, our DSFDA method addresses this challenge by employing a disentanglement process that effectively separates expression information from other facial identity attributes, resulting in a higher overall performance.
@@ 3.1 Notation
In the proposed method, $\mathbf{I}_{\text{id}}$ denotes the identity image with its corresponding one-hot label $y_{\text{id}}$, and $\mathbf{I}_{\text{exp}}$ denotes the expression image with its corresponding one-hot label $y_{\text{exp}}$. The generated image produced by the model is represented as $\mathbf{I}_{\text{g}}$, while $y_{\text{g}}$ is the one-hot label of the generated image $\mathbf{I}_{\text{g}}$. The number of identity classes is denoted as $c_{\text{id}}$, and the number of expression classes is represented as $c_{\text{exp}}$. The distribution of identity images is indicated by $p_{\text{id}}$, the distribution of expression images by $p_{\text{exp}}$, and the distribution of generated images by $p_{\text{g}}$. The reconstructed identity image is denoted as $\hat{\mathbf{I}}_{\text{id}}$, and the reconstructed expression image is represented as $\hat{\mathbf{I}}_{\text{exp}}$. The identity representation obtained from the identity encoder is indicated as $\mathbf{d}_{\text{id}}$, while $\mathbf{d}_{\text{g\_id}}$ refers to the identity representation of the generated image. Additionally, $n$ represents the number of training samples, and $\hat{y}_{i}^{j}$ is the $j$-th output value from the softmax classifier for the $i$-th sample.
3.2 Model Architecture
DSFDA is inspired by the architecture in [50], where expressions are disentangled from other facial information to improve expression classification performance. DSFDA relies on two encoders to disentangle identity and expression information from facial images for improved adaptation. It consists of a generator, with two encoders ($E_{\text{id}}$, $E_{\text{exp}}$) and a decoder ($D$), and two discriminators ($D_{\text{id}}$, $D_{\text{exp}}$) for identity and expression classification. The generator creates synthesized images the combining the identity of one image with the expression of another, and these images are verified by the discriminators. The model is trained using reconstruction, perceptual, and adversarial losses, along with an expression classifier ($C_{\text{exp}}$) trained on real data.
The training and adaptation architecture is a GAN model used originally to disentangle identity and expression from identity. Our architecture is thus made of different components: one generator, $G$, and two discriminators, $D_{\text{id}}$ and $D_{\text{exp}}$, respectively, for identity and expression. The generator is composed of one decoder ($D$) and two encoders; the first one specializes in extracting identity features, $E_{\text{id}}$, and the second one specializes in extracting expression features $E_{\text{exp}}$. A pair of images feed them $\mathbf{I}_{\text{id}}$ and $\mathbf{I}_{\text{exp}}$ from which a fake image $\mathbf{I}_{\text{g}}$ is generated with the identity of $\mathbf{I}_{\text{id}}$ and the pain level of $\mathbf{I}_{\text{exp}}$. Both discriminators $D_{\text{id}}$ and $D_{\text{exp}}$ are fed with this fake image to ensure that the correct identity and the correct expression have been generated. The generator is then once again used to generate, from $\mathbf{I}_{\text{g}}$ and the two real images $\mathbf{I}_{\text{id}}$ and $\mathbf{I}_{\text{exp}}$, two other fake images $\hat{\mathbf{I}}_{\text{id}}$ and $\hat{\mathbf{I}}_{\text{exp}}$, which are respectively the copies of $\mathbf{I}_{\text{id}}$ and $\mathbf{I}_{\text{exp}}$.
3.3 Objective Function

### v4

@@ Abstract
Facial Expression Recognition (FER) from videos is a crucial task in various application areas, such as human-computer interaction and health diagnosis and monitoring (e.g., assessing pain and depression). Beyond the challenges of recognizing subtle emotional or health states, the effectiveness of deep FER models is often hindered by the considerable inter-subject variability in expressions. Source-free (unsupervised) domain adaptation (SFDA) methods may be employed to adapt a pre-trained source model using only unlabeled target domain data, thereby avoiding data privacy, storage, and transmission issues. Typically, SFDA methods adapt to a target domain dataset corresponding to an entire population and assume it includes data from all recognition classes. However, collecting such comprehensive target data can be difficult or even impossible for FER in healthcare applications. In many real-world scenarios, it may be feasible to collect a short neutral control video (which displays only neutral expressions) from target subjects before deployment. These videos can be used to adapt a model to better handle the variability of expressions among subjects. This paper introduces the Disentangled SFDA (DSFDA) method to address the challenge posed by adapting models with missing target expression data. DSFDA leverages data from a neutral target control video for end-to-end generation and adaptation of target data with missing non-neutral data. Our method learns to disentangle features related to expressions and identity while generating the missing non-neutral expression data for the target subject, thereby enhancing model accuracy. Additionally, our self-supervision strategy improves model adaptation by reconstructing target images that maintain the same identity and source expression. Experimental results111 https://github.com/MasoumehSharafi/DSFDA/ on the challenging BioVid, UNBC-McMaster and StressID datasets indicate that our DSFDA approach can outperform state-of-the-art adaptation methods.
I INTRODUCTION
@@
Facial expression recognition (FER) has attracted much attention in computer vision due to the breadth of its applications, such as pain and depression assessment, stress estimation, and driver fatigue monitoring. Although state-of-the-art deep FER models [1, 2, 3, 4] have made significant progress, developing applications and models capable of characterizing expression remains a challenge. Deep learning models [5], [6], [7] have shown promising performance on various FER tasks. However, their performance can degrade significantly when there is a shift in data distributions between the source (lab) and target (operational) domains. Various unsupervised adaptation (UDA) methods have been
Figure 1: Illustration of SFDA settings using pain estimation as an example. (a) A standard subject-based SFDA setting where the source pre-trained model $F_{S}$ is adapted through a pseudo-labeling approach using unlabeled target video data from all classes (neutral and pain levels). (b) Our proposed subject-based disentangled SFDA (DSFDA) setting with neutral control target data. The generator $G$ is employed to synthesize the missing classes in the target domain based on the neutral control video. (c) Comparison of accuracy on BioVid data [8] of a source-only pre-trained Inception-V3 CNN ($F_{s}$ before adaptation), after adaptation with only neutral control data, and with our DSFDA method on neutral and generated pain data. (Lock symbols denote that labeled source data is unavailable during adaptation.)
proposed to address the distribution shifts between the source and the target domain [9, 10, 11, 12]. For instance, the work by Zeeshan et al. [11] introduces a subject-based multi-source domain adaptation (MSDA) method specifically tailored for FER tasks, where each subject is considered as an individual domain for fine-grained adaptation to the variability of expressions among subjects.
@@
State-of-the-art UDA methods [9, 12, 13, 14, 15] rely on the availability of the labeled source data during adaptation, which is impractical in real-world scenarios due to concerns for data privacy, data storage, transmission cost, and computation burden. Therefore, to reduce the reliance on source data during the adaptation process, several SFDA methods have been proposed [16, 17, 18, 19, 20].
@@
SFDA methods typically focus on the scenario where unlabeled target data from all classes is available for the adaptation process (see Fig. 1(a)). However, in real-world video-based health applications, person-specific data representing non-neutral expressions is typically costly or unavailable. A short neutral control video may, however, be collected for target individuals and used to personalize a model to the variability of an individual’s diverse expressions. In practice, collecting neutral target data for adaptation is generally easier than gathering non-neutral emotional data. Annotating neutral expressions is more straightforward and less subjective than interpreting emotional expressions, which can vary considerably among expert annotators. Creating scenarios to elicit specific emotions can be complex and time-consuming, and may raise ethical and privacy concerns. In contrast, capturing neutral expressions is less intrusive and may encourage more willing participation, as it avoids the need to act out emotional states.
@@
In this paper, a new personalized SFDA method is proposed for FER with our new setting (see Fig. 1(b)). The Disentangled SFDA (DSFDA) method is proposed to address the challenges of personalized SFDA using neutral control target data by generating the missing non-neutral expressive image data for target subjects. Using a dual-generator architecture pre-trained on source subjects, a neutral image is transformed into one that exhibits the missing non-neutral expression. By combining features from two input images, the method synthesizes expressive images that belong to the missing non-neural target expressions, enabling accurate FER predictions for target subjects even without explicit data for those expressions. Our approach is further enhanced using self-supervised learning, which ensures the reconstruction of input target images with the same identity and source expression. This process improves the model’s adaptability and predictive accuracy in the absence of source data. As shown in Fig. 1(c), DSFDA overcomes the challenges of domain shifts and incomplete target data, ensuring a high level of performance across diverse target domains.
Our contributions can be summarized as follows. (1) A novel SFDA setting is introduced, where all classes in the source data are utilized to pre-train the source model. For target adaptation, however, only neutral control target data is available to adapt the model for a target subject. (2) We propose the DSFDA method, a one-stage approach tailored to our novel setting, that combines data generation and target domain adaptation to simultaneously generate non-neutral target image data and adapt a FER model. (3) An extensive set of experimental results indicating that a FER model adapted using our DSFDA method can outperform state-of-the-art SFDA methods on the BioVid, UNBC-McMaster, and StressID datasets. The method scales well to the number of subjects and is robust across various subjects, target domains, and datasets.
II Related Work
II-A Source-Free Domain Adaptation
@@
Model-based methods: Conti et al. [29] introduced a cluster purity score to obtain reliable pseudo-labels using $k$-means clustering. The target feature extractor is then trained using two augmentations of input images instead of raw input data. The reduced target dataset is trained using the reliable pseudo-labels obtained with the self-supervised feature extractor. Guo et al. [30] incorporate Tsallis entropy into the information maximization constraint to enhance the confidence of prediction labels while reducing confidence bias.
@@
Unlike the promising results of SFDA methods in benchmark classification tasks, current SFDA approaches are typically designed under the assumption of an equal number of classes in both the source and target domains. However, in FER tasks, non-neutral expression data (missing classes) may be unavailable, with only a short neutral video available for adaptation. This leads to a label shift between the target adaptation and testing phases, ultimately degrading classification performance. To mitigate this shift between the source and target domains, generative models are utilized to synthesize the missing classes within the target domain. These models aim to produce high-quality images or videos that accurately represent the emotional expressions absent in the original target dataset. In the literature, a variety of techniques have been explored for this purpose, ranging from traditional generative adversarial networks (GANs) and variational autoencoders (VAEs) to more advanced frameworks such as conditional GANs, diffusion models, and disentanglement methods.
II-B Facial Expression Generation and Disentanglement
@@
Generative modeling, a branch of machine learning (ML), concentrates on understanding the underlying distribution of training data to enable the generation of new samples that mimic the statistical features of that data. Over time, significant progress has been made in this field. Most common techniques like GANs [31, 32, 33, 34, 35], VAEs [36, 37, 38], and diffusion models [39, 40, 41, 42] have significantly improved the quality and variety of the samples generated. For facial expression generation, [43] developed FExGAN-Meta for meta-humans, while ImaGINator [44] employed CGANs for creating facial expression video sequences. Niinuma et al. [45] used a GAN-based approach to synthesize facial expressions for AU intensity estimation, improving performance on imbalanced and domain-shifted datasets through personalized networks. Recent works have emphasized domain generalization and adaptation for cross-corpus facial expression recognition. Yin et al. [46] introduced the FATE model that uses contrastive learning and facial animation warping to achieve superior recognition of emotions in the cross-domain. Furthermore, DeepFN [47] proposed deep face normalization for robust recognition of action units between individuals. Kim et al. [48] leveraged single-image temporal cues for action unit detection, enhancing the understanding of subtle expressions.
In ML, disentanglement refers to the process of learning representations where distinct factors of variation in the data are captured in separate, interpretable components [49]. This property has proven beneficial in applications such as data augmentation, image synthesis, and improving model generalization. Several disentanglement methods have been proposed in the literature, and Wang et al. [50] suggest classifying these techniques according to various criteria. Evaluating disentanglement, however, remains a challenging task, as there is no universal metric capturing all aspects of factor separation. A recent survey provides a detailed review of existing disentanglement metrics, highlighting their limitations and suitability across use cases [51]. Disentanglement is particularly relevant in facial expression recognition (FER), where expression features are often entangled with identity-specific traits. Generative adversarial networks (GANs) are frequently employed to decouple expression and identity features, thereby enhancing both model performance and interpretability [7, 52, 53]. For instance, the IPD-FER [54] method disentangles identity, pose, and expression components, while Halawa et al. [55] propose a GAN-based approach that separates expression features without requiring identity labels.
@@
Beyond GANs, other disentanglement techniques include VAE methods [56]. Li et al. [57] use the Hilbert-Schmidt independence criterion to separate expression from identity features, showcasing mathematical criteria in disentanglement. Mutual information minimization is another strategy for this goal [58]. These studies illustrate various methods to achieve robust and interpretable disentangled representations in facial expression recognition.
Given only neutral videos in the target domain, we generate non-neutral expressions to adapt a source model. This two-stage approach re-frames the SFDA task by using a generator (e.g., VAE, GAN, or DM) trained on the source domain to synthesize expressive frames. These are combined with neutral target frames to adapt the model. However, current generators struggle to capture subtle pain intensities as a result of the complexity of nuanced expressions.
SOTA SFDA methods assume consistent classes across adaptation and testing, but FER tasks often lack non-neutral target expressions due to data collection challenges. This leads to label and domain shifts that affect performance. Generative models try to fill the gap but struggle with subtle expressions such as pain. Our DSFDA method overcomes this by disentangling expression from identity, increasing recognition accuracy.
@@ III Proposed Method
Notation. In the proposed method, $\mathbf{I}_{\text{t}}$ represents the target image, which can either be an expression image $\mathbf{I}_{\text{exp}}$ with its corresponding one-hot label $y_{\text{exp}}$ or an identity image $\mathbf{I}_{\text{id}}$ with its corresponding one-hot label $y_{\text{id}}$. The image generated by the model is represented as $\mathbf{I}_{\text{g}}$, while $y_{\text{g}}$ is the one-hot label of the generated image. The number of identity classes is denoted as $c_{\text{id}}$, and the number of expression classes is represented as $c_{\text{exp}}$. The distribution of identity images is indicated by $p_{\text{id}}$, the distribution of expression images by $p_{\text{exp}}$, and the distribution of generated images by $p_{\text{g}}$. The reconstructed target image is denoted as $\hat{\mathbf{I}}_{\text{t}}$, may represent either the reconstructed identity image $\hat{\mathbf{I}}_{\text{id}}$ or the reconstructed expression image $\hat{\mathbf{I}}_{\text{exp}}$. The identity representation extracted by the identity encoder is denoted as $\mathbf{d}_{\text{id}}$, while $\mathbf{d}_{\text{exp}}$ represents the expression embedding of the expression image. The source domain prototypes are indicated by $\mathbf{d}_{\text{s}}$. Additionally, $n$ represents the number of training samples, and $\hat{y}_{i}^{j}$ is the $j$-th output value from the softmax classifier for the $i$-th sample.
III-A Model Architecture
DSFDA builds upon the architecture proposed in [7], where facial expression information is disentangled from identity to enhance expression classification and domain adaptation. The model consists of a generator and two discriminators: the generator includes two encoders—an identity encoder ($E_{\text{id}}$) and an expression encoder ($E_{\text{exp}}$)—as well as a shared decoder ($D$). The discriminators, $D_{\text{id}}$ and $D_{\text{exp}}$, are constructed using the respective encoders followed by fully connected layers to perform identity and expression classification.
Given a pair of input images, $\mathbf{I}_{\text{id}}$ and $\mathbf{I}_{\text{exp}}$, the generator synthesizes an image $\mathbf{I}_{\text{g}}$ that combines the identity from $\mathbf{I}_{\text{id}}$ and the expression from $\mathbf{I}_{\text{exp}}$. This generated image is then assessed by the two discriminators to verify the preservation of the intended identity and expression. To ensure faithful disentanglement and reconstruction, the generator is further used to regenerate $\hat{\mathbf{I}}_{\text{id}}$ and $\hat{\mathbf{I}}_{\text{exp}}$ from $\mathbf{I}_{\text{g}}$, aiming to recover the original identity and expression inputs. Training is guided by a combination of reconstruction, perceptual, and adversarial losses, along with an expression classifier $C_{\text{exp}}$ trained on real data to support accurate expression recognition.
III-B Objective Functions